OpenStudy (anonymous):

solve and determine any extraneous solutions. 2*(sqrt4-2x)=2-x

OpenStudy (jack1):

\[\large 2 \times (\sqrt4-2x)=2-x\] or \[\large 2 \times (\sqrt{4-2x})=2-x\] ??? @berrypicking

OpenStudy (anonymous):

-6, 2 = x? considering is (1-(x/2))^2 = (4-2x)

OpenStudy (anonymous):

2nd equation 4-2x under radical sign

OpenStudy (anonymous):

yes the answers are -6 and 2 but not sure how to set up problem. I know I have to square both sides but don't know what to do with the 2 in front of the radical sign

OpenStudy (jack1):

\[\large 2 \times (\sqrt{4-2x})=2-x\] \[\large \sqrt{4-2x}=\frac {2-x}2\] \[\large (\sqrt{4-2x})^2=(\frac {2-x}2)^2\] \[\large 4-2x=\frac {x^2}4 -x +1\] \[\large 3-x=\frac {x^2}4 \] \[\large 12-4x={x^2} \] \[\large 12-4x-x^2=0 \] solve using quadratic method or factorize

OpenStudy (jack1):

as per previous answers: -6 or 2

OpenStudy (anonymous):

how did you come up with x^2/4

OpenStudy (jack1):

sorry dude, went to sleep \[\Large (\frac {2-x}2)^2\] \[\Large (\frac 22 - \frac x2)^2\] \[\Large (1 - \frac x2)^2\] \[\Large (1 - \frac x2) (1 - \frac x2)\] now use FOIL \[\Large (1*1 - 1*\frac x2 - 1*\frac x2 + (\frac x2)*(\frac x2)\] \[\Large =\frac {x^2}4 -x +1\]